# A Survey of Long-Document Retrieval in the PLM and LLM Era

### v1

@@ 4. PLM and LLM Era Models for Long-Document Retrieval
To systematically organize these cutting-edge approaches, this chapter categorizes them into several core paradigms, whose architectural overviews are illustrated in Fig. 3: (1) the Holistic Paradigm (detailed in Section 4.1), which aims to model and interact with the entire document as a single, indivisible unit; (2) the Divide-and-Conquer Paradigm (detailed in Section 4.2), which segments long documents into smaller pieces for localized processing before aggregating global information; (3) Long-Query Retrieval (detailed in Section 4.4), which tackles scenarios where queries themselves are long documents (e.g., legal case retrieval, patent prior-art search).Additionally, we discuss the Indexing-Structure-Oriented Paradigm in Section 4.3 — this paradigm focuses on innovating how documents are chunked and indexed (note: its workflows are highly diverse and difficult to consolidate into a single generalized diagram, so it is elaborated separately).
@@ 4.1.1. Naive Truncation Baselines
Early dense retrieval models such as DPR (Karpukhin et al., 2020) and ANCE (Xiong et al., 2020) highlighted the potential of pre-trained language models to surpass lexical retrieval by leveraging contrastive learning. However, they are constrained by BERT’s 512-token limit, and thus apply a naive truncation strategy—retaining only the first 512 tokens of a long document, based on the common assumption that important summary information resides at the beginning. This design introduces three critical flaws: (i) information loss, as relevant evidence beyond the cutoff is discarded; (ii) destruction of discourse structure, since long-range dependencies are severed; and (iii) positional bias, as the beginning of a document is always over-emphasized. Consequently, While DPR and ANCE pioneered dense retrieval, their truncation assumptions clearly fail in the context of long document IR. Truncation not only results in systematic information loss but also severely impairs the model’s ability to model discourse-level relevance. Furthermore, the performance ceiling of this approach is very low, making it difficult to serve as an effective baseline for long-document retrieval.
@@ 4.1.3. LLMs as Holistic Rerankers
With the advent of large language models , holistic modeling has advanced from efficient Transformers to end-to-end ranking. RankGPT (Sun et al., 2023) demonstrated that GPT-4 can act as a universal reranker by generating explicit ranked lists (e.g., “[2]¿[3]¿[1]”), rather than independent scores. Using a sliding-window strategy, it successfully reranked hundreds of passages, and permutation distillation further transferred this ability to a lightweight 440M DeBERTa, achieving superior nDCG@10 on BEIR with only 1k labels. RankGPT demonstrates the unique capabilities of LLM for listwise ranking. However, its reliance on sliding windows to join long documents can easily introduce query drift: the model may make inconsistent judgments for the same query in different windows. Furthermore, its high inference cost and prompt sensitivity make it difficult to deploy at scale.
@@ 4.1.5. Limitations and Open Problems
Holistic paradigms, while promising, remain constrained by efficiency, multimodality, and evaluation challenges. First, trade-offs between effectiveness and computational cost remain acute: million-token contexts (e.g., legal codes, technical manuals) remain impractical despite sparse attention and compression. Second, multimodal long documents—scientific papers with figures or news with embedded tables—remain largely unsupported; current VLMs (e.g., CLIP) lack robust alignment for 10k+ token multimodal contexts. Third, evaluation misalignment threatens progress: standard IR metrics (NDCG, Recall@k) fail to capture LLM-specific issues such as hallucinated evidence or failure to actually use context. Task-oriented metrics for faithfulness and attribution are urgently needed. Finally, robustness concerns—including adversarial injections, pretraining contamination, and domain adaptation fragility—underscore that LLMs cannot blindly replace IR pipelines.
@@ 4.2.2. Hierarchical Aggregation: PARADE, DRSCM, LTR-BERT, and MORES+
PARADE (Li et al., 2023b) encodes each $(q,p_{i})$ pair using a cross-encoder PLM to obtain passage representations $\mathbf{h}_{i}$, then aggregates them via multiple strategies, from simple pooling to learned networks. Among its learned strategies, PARADE-Attn uses a trainable vector to weigh passage importance, while the most powerful variant, PARADE-Transformer, employs a second Transformer encoder to model global dependencies between the passages. The final aggregated representation is then passed to a linear layer for scoring, allowing the model to form a holistic judgment based on evidence scattered across the text. The authors show that such learned aggregation methods consistently outperform simple heuristics, demonstrating that relevance in long documents often emerges from these cross-passage interactions.
@@
LTR-BERT (Wang et al., 2024) decouples heavy offline document encoding from lightweight online query processing. Long documents are segmented and encoded offline, producing compressed passage embeddings. At query time, a short query is expanded and encoded online, then matched to stored embeddings using a parameter-free late interaction mechanism. This lightweight matching operates token-wise: for each query token, it finds the most similar document token within a passage, and the final score is computed by comparing the averaged vectors of the query and these selected best-match tokens. Notably, the model is trained efficiently on short-text pairs and uses a BERT-MaxP-style final aggregation, but achieves its speed by replacing the expensive cross-encoder with its parameter-free calculation. The result is a large efficiency gain—processing up to 333 times more documents per millisecond—while still outperforming cross-encoder baselines on MS MARCO document ranking.
@@ 4.3. Indexing-Structure-Oriented Paradigm
RAPTOR(Sarthi et al., 2024) organizes a document into a semantic tree through a process of recursive clustering and summarization. As visualized in Figure 5 (c), this bottom-up construction begins with initial text chunks (leaf nodes). These chunks are then recursively clustered based on semantic similarity and summarized by an LLM to form higher-level nodes. The diagram illustrates this iterative process, showing how multiple layers of summaries are built until a single root node is formed, effectively grouping semantically related information regardless of its original position in the text. For querying, RAPTOR’s more robust ”collapsed tree” strategy was found to be superior to a simple top-down traversal. This approach works by flattening all nodes from the tree—both original chunks and all levels of summaries—into a single pool. A standard dense retrieval is then performed across this entire pool, allowing the query to match information at any level of abstraction. By leveraging these multiple levels of abstraction, RAPTOR achieves notable gains on complex question-answering tasks. For instance, on the QuALITY dataset(Pang et al., 2021b) , it boosts accuracy by over 8 percentage points compared to a DPR baseline, highlighting the potential of hierarchical, abstraction-aware indexing.
@@ 4.4. Long-Query Retrieval: The Query-by-Document Task
To address these challenges, sentence-level bi-encoder models have been proposed, including one based on a Re-ranking with Proportional Relevance Score (RPRS(Askari et al., 2024). Built on SBERT, RPRS operates by segmenting both the query $Q$ and candidate document $D$ into sentences. Each sentence is then passed through an encoder $E(\cdot)$ to produce its vector embedding:
@@
In a modern retrieval pipeline, RPRS serves as a versatile and efficient re-ranker. By explicitly balancing query and document coverage, it avoids the pitfalls of methods that may over-emphasize a few salient passages while ignoring broad topical alignment. The source paper shows that while RPRS is robust, its effectiveness can be further improved by tuning its three parameters ($n,k1,b$) on a small amount of labeled data. Future work includes adapting the method for first-stage retrieval and exploring techniques to make it parameter-free by dynamically computing its parameters.
@@ 5.1. Datasets
A selection of commonly used long-document retrieval datasets is summarized in Table 2.
Table 2. Long Document Retrieval Dataset
Dataset Corpus size Queries / splits Avg. doc/para length Label / notes
TREC Robust04 528,155 documents 250 topics (Robust set) Avg. doc length:  1,000 words graded relevance; news/ad-hoc reranking. (Voorhees, 2004, 1996; Huston and Croft, 2014)
Gov2 / Terabyte Track 25,205,179 documents topic counts vary by year (e.g., many Terabyte tasks used 50 topics) avg. 17.7 KB / doc (collection size $\approx$426 GB) binary/pooled judgments; large-scale web crawl. (Büttcher et al., 2006)
ClueWeb12 733,019,372 pages (Cat. A) used with TREC Web tracks (topic sets vary) Collection size:  27.3 TB (uncompressed); Avg. doc length:  40 KB web-scale, noisy pages; used for robustness / web search experiments. (Lemur Project, 2012)
TREC Deep Learning (2019–2020) / MS MARCO (doc) $\sim$3.2M documents (doc task) train $\approx$367,013 queries; reusable doc test set: 43 queries avg. $\sim$1,000 tokens / document (typical) human labels (sparse/graded); de-facto neural reranking benchmark. (Craswell et al., 2020)
MLDR (2024) corpus total $\approx$493,709 documents (13 langs) Queries (total over 13 langs): Train 41,434, Dev 2,600, Test 3,800 avg. doc length $\approx$4,737 (units as reported per-language) multilingual long-document retrieval; queries generated via LLM prompts. (Chen et al., 2024)
LONGEMBED task-wise corpora e.g. NarrativeQA: 10,449 queries / 355 docs; QMSum: 1,527 / 197; 2WikiMultihopQA: 300 / 300; SummScreenFD: 336 / 336 e.g., NarrativeQA avg. 50,474 words/doc; QMSum avg. 10,058 words/doc; (also synthetic tasks with contexts up to 32k tokens) benchmark for long-context retrieval; mixture of long real-world and synthetic tasks. (Zhu et al., 2024)
Multi-CPR (Chinese) passages per-domain: E-comm 1,002,822; Ent-video 1,000,000; Medical 959,526 train (per domain) $\approx$100,000; test 1,000 (per domain) avg. passage length: E-comm 32.96 words; Ent-video 27.45; Medical 122.02 large multi-domain Chinese passage retrieval with human annotations. (Long et al., 2022)
TREC Robust04. A classical ad-hoc news benchmark with graded relevance and strong lexical baselines (Voorhees, 2004). Still widely used for document reranking and for studying passage aggregation. Use the maintained distributions (disks45/nocr/trec-robust-2004) for reproducibility (Voorhees, 1996; Huston and Croft, 2014).
Gov2 / Terabyte Track. Large-scale web crawl used in the Terabyte Tracks with binary judgments (Büttcher et al., 2006). Documents are multi-paragraph; depth of pooling and label sparsity require careful reporting of evaluation cutoffs.
ClueWeb12. Web-scale corpus (733M English pages) supporting ad-hoc and web search studies; commonly paired with TREC Web Tracks (Lemur Project, 2012). Long, noisy pages make it suitable for testing robustness and dispersed relevance.
TREC Deep Learning (DL) 2019–2020. Document ranking over MS MARCO web documents ($\sim$3.2M; average length $\sim$1K tokens) with graded labels (Craswell et al., 2020). While documents are medium-length, DL’19/’20 remains the de facto neural reranking benchmark.
MLDR (2024). A multilingual long-document retrieval suite (13 languages) with longer articles and generated queries/answers (Chen et al., 2024). Useful for testing cross-lingual generalization under long contexts.
LONGEMBED. A benchmark(Zhu et al., 2024) developed to evaluate long-context retrieval capabilities.It includes synthetic tasks, such as personalized password retrieval and needle-in-a-haystack retrieval, along with real-world tasks like NarrativeQA, 2WikiMultihopQA, QMSum, and SummScreenFD, designed to measure the ability of embedding models to process long text designed to measure the ability of embedding models to process long texts.
Multi-CPR This is a multi-domain Chinese language dataset for paragraph retrieval(Long et al., 2022). The dataset is collected from three different domains: e-commerce, entertainment video, and healthcare. Each dataset contains millions of paragraphs and a certain number of manually annotated query-paragraph pairs.
@@ 5.2. Evaluation protocols and metrics
5.2.1. Traditional IR Metrics for long-document retrieval
This section systematically reviews the standard evaluation metrics widely used in traditional information retrieval (IR), including their mathematical definitions, computational logic, and their applicability and limitations in long-document retrieval (LDR) scenarios. These metrics are the cornerstone of the IR field, but they exhibit significant flaws when processing long-documents (such as legal contracts and scientific papers), primarily due to evidence dispersion and label sparsity. Traditional information retrieval metrics focus on relevance and efficiency, making them suitable for short document retrieval, but should be used with caution in LDR. The definition, formula, and calculation examples for each metric are as follows:
Let $Q$ be queries, $L_{q}=\{d_{1},\dots,d_{k}\}$ the ranked list for $q$, and $\mathrm{rel}_{i}\in\{0,\dots,G\}$ the (graded) relevance at rank $i$.
Precision@$k$: Measures the proportion of relevant documents in the top k search results. Reflects the accuracy of the system in top-k results.
(9) $\mathrm{P@}k(q)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}[\mathrm{rel}_{i}>0],$
where $\mathrm{rel}_{i}$ represents the relevance score (binary or hierarchical) of the $i$th document, and $I$ is an indicator function (1 if relevant, 0 otherwise). It is simple and intuitive, but is susceptible to label sparsity. In long documents, relevant evidence is scarce and scattered, resulting in large fluctuations in the P@k value (e.g., the variance on the TREC DL dataset is as high as ±10
Recall@$k$: The evaluation system is used to evaluate the proportion of relevant documents retrieved by the system to all relevant documents, emphasizing coverage.
@@ 5.2.1. Traditional IR Metrics for long-document retrieval
Where $\mathrm{R}_{q}$is the set of all relevant documents for query $\mathrm{q}$. Recall@k is more stable in long documents with scattered evidence, but ignores position information (deeply related documents may be underestimated) and relies on complete related annotations (which are often missing in practice).
MRR. Mean Reciprocal Rank (MRR): Calculates the mean of the reciprocal of the first relevant document ranking to highlight the importance of the top-ranked results.
@@
where $\mathrm{rank}_{q}$ is the first relevant rank. MRR emphasizes early relevance and is suitable for short query scenarios; however, in LDR, because relevant evidence may be scattered in the middle or end of the document (such as the results section of a scientific paper), MRR will systematically underestimate the effect.
nDCG@$k$. Normalized Discounted Cumulative Gain (nDCG@$k$) combines graded relevance and position discount to evaluate the cumulative gain of the top-k results.
(12) $\mathrm{DCG@}k(q)=\sum_{i=1}^{k}\frac{2^{\mathrm{rel}_{i}}-1}{\log_{2}(i+1)},$
(13) $\mathrm{nDCG@}k(q)=\mathrm{DCG@}k/\mathrm{IDCG@}k.$
where nDCG@$k$ is the ideal DCG (gain under perfect sorting). nDCG accounts for graded relevance, but in long-context settings, deep relevant passages can receive disproportionately low gain, potentially underestimating systems designed for exhaustive retrieval.In long-text retrieval, discounting deeply relevant passages can lead to overestimation of under-coverage systems;
MAP. Mean Average Precision (MAP): Calculates the average precision (AP) across all queries, emphasizing recall and ranking of relevant documents overall.
(14) $\mathrm{AP}(q)=\frac{1}{|R_{q}|}\sum_{i=1}^{N}\mathrm{P@}i(q)\cdot\mathbb{I}[\mathrm{rel}_{i}>0];$
(15) $\mathrm{MAP}=\frac{1}{|Q|}\sum_{q}\mathrm{AP}(q).$
MAP rewards retrieving all relevant units, yet its sensitivity to incomplete judgments makes it problematic for web-scale or long-document corpora with shallow pooling. Because it is highly sensitive to incomplete judgments, it is recommended to provide both ”judgment coverage” and ”incomplete judgment handling strategies” on datasets such as MS MARCO and TREC DL.
ERR. Expected Reciprocal Rank (ERR) models the process of a user sequentially examining results from the top, with a certain probability of being satisfied and stopping at each rank. the graded relevance label at rank $i$. ERRmodels the process of a user sequentially examining results from the top of the ranking, with a certain probability of being satisfied and stopping at each position.
(16) $R_{i}=\frac{2^{\mathrm{rel}_{i}}-1}{2^{G}},$
where $G$ denotes the maximum relevance grade. The probability that the user stops exactly at rank $i$ is given by
(17) $P_{i}=R_{i}\prod_{j=1}^{i-1}(1-R_{j}),$
which accounts for the chance that no higher-ranked document was satisfying and the current one is. The Expected Reciprocal Rank is then defined as
(18) $\mathrm{ERR}@K=\sum_{i=1}^{K}\frac{1}{i}\,P_{i}=\sum_{i=1}^{K}\frac{1}{i}\left(R_{i}\prod_{j=1}^{i-1}(1-R_{j})\right).$
Intuitively, ERR represents the expected reciprocal rank at which a user finds a satisfying document, thereby emphasizing the importance of placing highly relevant items early in the ranking. It is particularly suitable in evaluation scenarios with graded relevance judgments, such as web search or recommendation, where user attention is limited and users are likely to stop once they encounter a highly relevant result.
5.2.2. Efficiency-Utility Trade-offs
Efficiency metrics.
@@ 5.2.2. Efficiency-Utility Trade-offs
Beyond accuracy, efficiency is critical in LDR. Measures include end-to-end latency (per query), index size and GPU usage. These factors determine practical deployability and comparability across methods.
Robustness and generalization. Evaluate performance across domains (e.g., news, scientific papers, web) and under noisy conditions (OCR, HTML artifacts). Cross-domain benchmarks (e.g., BEIR (Thakur et al., 2021)) are recommended to measure robustness.
5.3. Benchmarking practices and common pitfalls
  • Label sparsity and pooling bias. On web-scale corpora (Gov2/ClueWeb12), shallow pooling yields few judged relevant documents per query, inflating variance. Always report the candidate generation recipe and depth.
  • Query realism. Many benchmarks rely on short keyword queries; where the application requires natural-language, multi-constraint questions, consider complementary long-query (QBD) or structured QA resources (Sec. 5.1).
  • Structure/layout neglect. Flattening long documents discards section boundaries, figures, and tables that guide retrieval. Prefer structure-aware datasets (WikiWeb2M-QA) or multimodal suites (MMDOCIR) when appropriate.
  • Unit mismatch. Using document-level metrics when supervision is segment-level can mask failure modes. Report both segment- and document-level performance when feasible.
  • Reproducibility. Fix random seeds; publish depth-$k$ cutoffs, negative sampling, reranking pool sizes, and efficiency (latency/memory) under the same hardware for fair comparisons.

### v2

@@ 4. PLM and LLM Era Models for Long-Document Retrieval
To systematically organize these cutting-edge approaches, this chapter categorizes them into several core paradigms, whose architectural overviews are illustrated in Fig. 3: (1) the Holistic Paradigm (detailed in Section 4.1), which aims to model and interact with the entire document as a single, indivisible unit; (2) the Divide-and-Conquer Paradigm (detailed in Section 4.2), which segments long documents into smaller pieces for localized processing before aggregating global information; (3) Long-Query Retrieval (detailed in Section 4.4), which tackles scenarios where queries themselves are long documents (e.g., legal case retrieval, patent prior-art search). Additionally, we discuss the Indexing-Structure-Oriented Paradigm in Section 4.3, this paradigm focuses on innovating how documents are chunked and indexed (note: its workflows are highly diverse and difficult to consolidate into a single generalized diagram, so it is elaborated separately).
@@ 4.1.1. Naive Truncation Baselines
Early dense retrieval models such as DPR (Karpukhin et al., 2020) and ANCE (Xiong et al., 2020) highlighted the potential of pre-trained language models to surpass lexical retrieval by leveraging contrastive learning. However, they are constrained by BERT’s 512-token limit, and thus apply a naive truncation strategy: retaining only the first 512 tokens of a long document, based on the common assumption that important summary information resides at the beginning. This design introduces three critical flaws: (i) information loss, as relevant evidence beyond the cutoff is discarded; (ii) destruction of discourse structure, since long-range dependencies are severed; and (iii) positional bias, as the beginning of a document is always over-emphasized. Consequently, While DPR and ANCE pioneered dense retrieval, their truncation assumptions clearly fail in the context of long document IR. Truncation not only results in systematic information loss but also severely impairs the model’s ability to model discourse-level relevance. Furthermore, the performance ceiling of this approach is very low, making it difficult to serve as an effective baseline for long-document retrieval.
@@ 4.1.3. LLMs as Holistic Rerankers
With the advent of large language models , holistic modeling has advanced from efficient Transformers to end-to-end ranking. RankGPT (Sun et al., 2023) demonstrated that GPT-4 can act as a universal reranker by generating explicit ranked lists (e.g., “[2]¿[3]¿[1]”), rather than independent scores. Using a sliding-window strategy, it successfully reranked hundreds of passages, and permutation distillation further transferred this ability to a lightweight 440M DeBERTa, achieving superior nDCG@10 on BEIR with only 1k labels. RankGPT demonstrates the unique capabilities of LLM for listwise ranking. However, its reliance on sliding windows to join long documents can easily introduce query drift: the model may make inconsistent judgments for the same query in different windows. Furthermore, its high inference cost and prompt sensitivity make it difficult to deploy at scale (Hagström et al., 2025).
@@ 4.1.5. Limitations and Open Problems
Holistic paradigms, while promising, remain constrained by efficiency, multimodality, and evaluation challenges. First, trade-offs between effectiveness and computational cost remain acute: million-token contexts (e.g., legal codes, technical manuals) remain impractical despite sparse attention and compression. Second, multimodal long documents: scientific papers with figures or news with embedded tables, remain largely unsupported; current VLMs (e.g., CLIP) lack robust alignment for 10k+ token multimodal contexts. Third, evaluation misalignment threatens progress: standard IR metrics (NDCG, Recall@k) fail to capture LLM-specific issues such as hallucinated evidence or failure to actually use context. Task-oriented metrics for faithfulness and attribution are urgently needed. Finally, robustness concerns including adversarial injections, pretraining contamination, and domain adaptation fragility underscore that LLMs cannot blindly replace IR pipelines.
@@ 4.2.2. Hierarchical Aggregation: PARADE, DRSCM, LTR-BERT, and MORES+
PARADE (Li et al., 2023b) encodes each $(q,p_{i})$ pair using a cross-encoder PLM to obtain passage representations $h_{i}$, then aggregates them via multiple strategies, from simple pooling to learned networks. Among its learned strategies, PARADE-Attn uses a trainable vector to weigh passage importance, while the most powerful variant, PARADE-Transformer, employs a second Transformer encoder to model global dependencies between the passages. The final aggregated representation is then passed to a linear layer for scoring, allowing the model to form a holistic judgment based on evidence scattered across the text. The authors show that such learned aggregation methods consistently outperform simple heuristics, demonstrating that relevance in long documents often emerges from these cross-passage interactions.
@@
LTR-BERT (Wang et al., 2024) decouples heavy offline document encoding from lightweight online query processing. Long documents are segmented and encoded offline, producing compressed passage embeddings. At query time, a short query is expanded and encoded online, then matched to stored embeddings using a parameter-free late interaction mechanism. This lightweight matching operates token-wise: for each query token, it finds the most similar document token within a passage, and the final score is computed by comparing the averaged vectors of the query and these selected best-match tokens. Notably, the model is trained efficiently on short-text pairs and uses a BERT-MaxP-style final aggregation, but achieves its speed by replacing the expensive cross-encoder with its parameter-free calculation. The result is a large efficiency gain: processing up to 333 times more documents per millisecond, while still outperforming cross-encoder baselines on MS MARCO document ranking.
@@ 4.2.3. Key Passage/Block Selection: IDCM, KeyB, KeyB2, and DCS
BReps (Li et al., 2025) introduces a block-level representation framework for LDR. Documents are segmented into short, semantically coherent blocks, each encoded independently by a LLM. At query time, the query embedding is matched against all block embeddings, and the top-$k$ relevant block scores are aggregated to produce the document-level score. This design enables offline pre-computation of block representations and lightweight online scoring, while capturing fine-grained matching signals that are often lost in single-vector representations. Compared with dense (e.g., RepLLaMA) and late-interaction baselines, BReps achieves higher effectiveness, demonstrating consistent gains across multiple long document benchmarks.
@@ 4.3. Indexing-Structure-Oriented Paradigm
RAPTOR(Sarthi et al., 2024) organizes a document into a semantic tree through a process of recursive clustering and summarization. As visualized in Figure 5 (c), this bottom-up construction begins with initial text chunks (leaf nodes). These chunks are then recursively clustered based on semantic similarity and summarized by an LLM to form higher-level nodes. The diagram illustrates this iterative process, showing how multiple layers of summaries are built until a single root node is formed, effectively grouping semantically related information regardless of its original position in the text. For querying, RAPTOR’s more robust ”collapsed tree” strategy was found to be superior to a simple top-down traversal. This approach works by flattening all nodes from the tree (both original chunks and all levels of summaries) into a single pool. A standard dense retrieval is then performed across this entire pool, allowing the query to match information at any level of abstraction. By leveraging these multiple levels of abstraction, RAPTOR achieves notable gains on complex question-answering tasks. For instance, on the QuALITY dataset(Pang et al., 2021b) , it boosts accuracy by over 8 percentage points compared to a DPR baseline, highlighting the potential of hierarchical, abstraction-aware indexing.
@@ 4.4. Long-Query Retrieval: The Query-by-Document Task
To address these challenges, sentence-level bi-encoder models have been proposed, including one based on a Re-ranking with Proportional Relevance Score (RPRS) (Askari et al., 2024). Built on SBERT, RPRS operates by segmenting both the query $Q$ and candidate document $D$ into sentences. Each sentence is then passed through an encoder $E(\cdot)$ to produce its vector embedding:
@@
In a modern retrieval pipeline, RPRS serves as a versatile and efficient re-ranker. By explicitly balancing query and document coverage, it avoids the pitfalls of methods that may over-emphasize a few salient passages while ignoring broad topical alignment. The source paper shows that while RPRS is robust, its effectiveness can be further improved by tuning its three parameters ($n,k_{1},b$) on a small amount of labeled data. Future work includes adapting the method for first-stage retrieval and exploring techniques to make it parameter-free by dynamically computing its parameters.
@@ 5.1. Datasets
TREC DL Document Ranking Track (2019–2023). The document ranking task in the TREC Deep Learning (DL) track has been evaluated annually since 2019, with two subtasks: (i) full ranking from the entire corpus and (ii) top-100 re-ranking using organizer-provided candidates. Graded relevance labels (0–3) are used for testing, and NDCG@10 is the primary metric.
  • MS MARCO v1 (2019–2020): The 2019 and 2020 editions used the MS MARCO v1 document collection of approximately 3.21M web documents. For training, document labels were transferred from passage labels—a document was marked relevant if it contained a relevant passage. For evaluation, however, NIST provided independent human judgments at the document level (43 topics in 2019; 45 topics in 2020). Both full ranking and top-100 re-ranking subtasks were offered.
  • MS MARCO v2 (2021–2023): From 2021, the track switched to the larger MS MARCO v2 corpus with about 11.9M documents. In 2021, document evaluation combined independent document judgments with labels propagated from passage annotations (57 topics). In 2022 and 2023, evaluation resources focused on building a reusable passage test set; the document task’s labels were inferred from passage judgments (76 topics in 2022; 82 topics in 2023).
TREC Robust04. The TREC Robust04111https://trec.nist.gov/data/robust/04.guidelines.html dataset is a classic benchmark in the information retrieval field, widely used in research on long document retrieval tasks (Voorhees, 2004). This dataset originates from the TREC Robust Track 2004, with the core goal of evaluating the robustness of retrieval systems when facing difficult queries. The TREC Robust document collection is from TREC disks 4 and 5. It contains approximately 528,000 long news documents covering multiple news sources, 250 topics, and approximately 31,000 qrels. Relevance judgments in Robust04 are binary: documents are either marked as relevant (1) or non-relevant (0).
GOV2 / Terabyte Track. Gov2222https://ir.dcs.gla.ac.uk/test_collections/gov2-summary.htm is a large-scale web page corpus released by TREC in 2004 and is the core dataset of the TREC Terabyte Track (2004–2006) (Büttcher et al., 2006). GOV2 contains documents resulting from a crawl of .gov websites made in early 2004 and contains approximately 25 million web pages, with documents truncated to 256 kilobytes (KB), for an average document size of approximately 17.7 KB. It contains 150 queries with three level judgment: 0 (irrelevant), 1 (relevant) or 2 (very relevant).
ClueWeb12. The ClueWeb12 dataset is designed to support research in information retrieval and related human language technologies. It contains 733,019,372 English web pages (Lemur Project, 2012). The dataset is divided into segments, each containing approximately 50 million documents. These documents are HTML web pages containing noisy content such as navigation, advertisements, and templates. Long, noisy pages make it suitable for testing robustness and dispersed relevance.
MQ2007/2008. The MQ2007 and MQ2008 datasets are part of the LETOR 4.0 benchmark (Qin and Liu, 2013), constructed from the GOV2 web collection of approximately 25 million documents and the TREC Million Query Tracks of 2007 and 2008. MQ2007 contains around 1,700 queries with roughly 65K judged query-document pairs, while MQ2008 includes about 800 queries with 14K pairs. Each query-document pair is represented by a 46-dimensional feature vector, incorporating various retrieval signals such as BM25 scores, language model features, and link-based features. Relevance annotations are graded on three levels (0, 1, 2), and the benchmark provides 5-fold cross-validation partitions for training, validation, and testing. Although originally designed for learning-to-rank research, MQ2007/2008 are also cited in the context of long document retrieval because they are derived from full web pages in the GOV2 corpus. The relevance judgment of query-document labels enable training and evaluation of long document IR models.
MLDR. MLDR is a multilingual long document retrieval dataset based on Wikipedia, Wudao, and mC4, covering 13 typologically diverse languages (Chen et al., 2024). It is constructed by extracting long articles from these datasets and randomly selecting paragraphs from them. GPT-3.5 is then used to generate questions based on these paragraphs. The generated questions and the extracted articles constitute the new text pairs in the dataset. It contains 41,434 training documents, 2,600 test documents, and 3,800 validation documents, with an average document length of 4,737.
@@ 5.2. Evaluation protocols and metrics
This section reviews standard IR metrics that are also widely used for LDR. We first present their definitions, followed by a summary of potential challenges and considerations specific to LDR scenarios.
Let $Q$ be the set of queries, $L_{q}=\{d_{1},\dots,d_{N}\}$ the ranked list for query $q$, and $\mathrm{rel}_{i}\in\{0,\dots,G\}$ the (graded) relevance of $d_{i}$ at rank $i$. Let $\mathbb{I}[\cdot]$ be the indicator function and $R_{q}=\{d:\mathrm{rel}(q,d)>0\}$ the set of relevant documents for $q$.
Precision@$k$.
(9) $\mathrm{P@}k(q)=\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}[\mathrm{rel}_{i}>0].$
Measures the proportion of relevant documents in the top-$k$ results.
Recall@$k$.
@@
Evaluates coverage of relevant documents retrieved within the top-$k$.
Mean Reciprocal Rank (MRR).
@@
where $\mathrm{rank}_{q}$ is the position of the first relevant document. Captures how early the first relevant document appears.
Normalized Discounted Cumulative Gain (nDCG@$k$).
(12) $\mathrm{DCG@}k(q)=\sum_{i=1}^{k}\frac{2^{\mathrm{rel}_{i}}-1}{\log_{2}(i+1)},\qquad\mathrm{nDCG@}k(q)=\frac{\mathrm{DCG@}k(q)}{\mathrm{IDCG@}k(q)}.$
Accounts for graded relevance and rank position through logarithmic discounting.
Mean Average Precision (MAP).
(13) $\mathrm{AP}(q)=\frac{1}{|R_{q}|}\sum_{i=1}^{N}\mathrm{P@}i(q)\cdot\mathbb{I}[\mathrm{rel}_{i}>0],\qquad\mathrm{MAP}=\frac{1}{|Q|}\sum_{q}\mathrm{AP}(q).$
Aggregates precision across recall levels to evaluate overall ranking quality.
@@
Beyond accuracy, efficiency is also critical in LDR. Measures include end-to-end latency (per query), index size and GPU usage. These factors determine practical deployability and comparability across methods.
Discussion. Traditional IR metrics remain the foundation of evaluation in LDR, but several characteristics of long documents complicate their interpretation. First, judgments are typically applied at the document level, which obscures finer distinctions such as “partially relevant” or “locally relevant” segments. Second, relevance evidence can be dispersed across different sections, while these metrics only consider whether the document as a whole is relevant. Third, coarse relevance scales are often insufficient to reflect nuanced utility in long documents. Future work may extend metrics to incorporate evidence localization and finer-grained relevance levels.